Brittney is 5 years older than her brother Brandon. 3 years from now the sum of their age will be 23. How old is Brandon?
Answer by macston(5194)

.
The current sum of their ages is (23 yrs- 6yrs)=17 yrs
B=Brandon's age; B+5yrs=Brittney's age
B+B+5yrs=17yrs
2B=12yrs
B=6yrs.
ANSWER: Brandon is 6 years old.
.
CHECK:
Brittney=B+5=6+5=11 years old
In three years:
(11yrs+3yrs)+(6yrs+3yrs)=23yrs
24yrs+9yrs=23yrs
23yrs=23yrs